a collimated laser beam with a divergence of 4.1 milliradians is incident on a lens of focal length 250mm. Calculate the spot size of laser light at the focal point of the lens
Answer #1

Using the equation we have

\theta = 1.22*\lambda / D

4.1*10^-3 = 1.22*250*10^-3 / D

D = 74.4 m

so spot size = D/2 = 74.4 /2 37.2 m